If Log of 32.5_to the base 10 is 1.5118834 than the number whose log to the
same base is 3.5118834 is
a) 0.00325
b) 0.000325
c) 0.0000325
d) none

Answer by stanbon(75887)

If Log of 32.5_to the base 10 is 1.5118834 than the number whose log to the
same base is 3.5118834 is
a) 0.00325
b) 0.000325
c) 0.0000325
d) none
-----------------
Log32.5 = 1.5118814 means the following:
10^1.5118814 = 32.5
----------
So 10^3.5118814= 10^2*10^1.5118814= 100(32.5)=3250
Therefore log3250=3.5118834
=========
Cheers,
Stan H.